Three charges Q1= 2 microcoulomb (0,0)cm, Q2= 4 microcoulomb (0,10)cm and Q3= -3 microcoulomb (15,0)cm acts on one number another. Find the magnitude of the total electric force and the angle that the total electric force makes with the positive x-axis


1
Expert's answer
2021-05-25T10:11:40-0400

"Q_1=2\times10^{-6}\space C"

"Q_2=4\times10^{-6}\space C"

"Q_3=-3\times10^{-6}\space C"


"F_{13}=k\dfrac{Q_1Q_3}{(15\times10^{-2})^2}=9\times10^9\dfrac{(2\times10^{-6})(3\times10^{-6})}{(15\times10^{-2})^2}"

"F_{13}=2.4\space N"


"F_{23}=k\dfrac{Q_2Q_3}{(18.02\times10^{-2})^2}=9\times10^9\dfrac{(4\times10^{-6})(3\times10^{-6})}{(18.02\times10^{-2})^2}"

"F_{23}=3.32\space N"


Angle between "F_{23}" and "F_{13}"

"\tan\theta=\dfrac{10}{15}"

"\theta=33.69\degree"


Resultant force,

"F=\sqrt{F_{13}^2+F_{23}^2-2F_{13}F_{23}\cos\theta}"

"F=1.876\space N"


Angle made by resultant force with x axis

"\phi=180\degree-\dfrac{\theta}{2}=163.155\degree"